Company Lagrange has three close peers whose current stock prices per share (P) and NTM EPS values are as follows. Peer 1: P = $60, EPS = $7; Peer 2: P = $70, EPS = $8; Peer 3: P = $80, EPS = $9. Lagrange's NTM EPS = $10. What should be its share price based on the maximum PE multiple of its three close peers?
Answer #1

PE Ratio of Peer 1 = 60/7 = 8.57

PE Ratio of Peer 2 = 70/8 = 8.75

PE Ratio of Peer 3 = 80/9 = 8.89

EPS of NTM = $10

Price of NTM = 10(8.89)

Price of NTM = $88.90